Please help me with the following word problem .A scoreboard is 10 meters squared,the scoreboard was enlarged, the length was increased by 1m and the width by 3 meters, by doing so the area of the scoreboard tripled. determine the original length and width
Answer by rothauserc(4718)

Area of the scoreboard is 10 m^2 and we are given,
l*w = 10
(l+1)*(w+3) = 30
solve first formula for l, l = 10/w
substitute for l in the second equation
(10/w +1)*(w+3) = 30
((10+w)/w) * (w+3) = 30
multiply both sides of = by w
(w+10) * (w+3) = 30w
w^2 +13w +30 = 30w
w^2 -17w +30 = 0
factor equation
(w-15)*(w-2) = 0
w = 15 and w=2
we have two possibilities
w = 2 m, and l = 5 m and
w = 15 m and l = 2/3 m